# A short proof of Isserlis’ theorem

### v1

@@ 1. Isserlis’s theorem
Let ${\bf Y}=(Y^{1},Y^{2},\ldots,Y^{n})^{T}\sim{\mathcal{N}}_{n}({\bf 0},\Sigma)$ be a multivariate normally distributed vector in ${\mathbb{R}}^{n}$ with zero mean. Then
(1) ${\mathbb{E}}(Y^{1}Y^{2}\cdots Y^{n})=\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathbb{E}}(Y^{\ell}Y^{r}).$
The sum is over all partitions of $n$ into disjoint pairs, denoted $\operatorname{PP}(n)$. For odd $n$, $\operatorname{PP}(n)=\emptyset$ is empty and (1) reduces to $E(Y^{1}\cdot Y^{2}\cdots Y^{n})=0$, and we recover that odd moments of normal centred distributions always vanish. For even $n$ with $n=2k$, the number of elements of $\operatorname{PP}(n)$ is exactly #$\#\operatorname{PP}(n)=\frac{(2k)!}{2^{k}k!}.$ For $n=4$, we have #$\#\operatorname{PP}(4)=3$ with
@@
${\mathbb{E}}(Y^{1}Y^{2}Y^{3}Y^{4})={\mathbb{E}}(Y^{1}Y^{2}){\mathbb{E}}(Y^{3}Y^{4})+{\mathbb{E}}(Y^{1}Y^{3}){\mathbb{E}}(Y^{2}Y^{4})+{\mathbb{E}}(Y^{1}Y^{4}){\mathbb{E}}(Y^{2}Y^{3}).$
@@ 2. Isserlis’ Theorem as a corollary of the invariant tensor theorem
We shall derive formula (1) as a consequence of the invariant tensor theorem for isotropic tensors [13, Sec. II.9][7, §33.2] stating that the linear space of isotropic tensors is spanned by certain elementary isotropic tensors with two inputs.
@@
The first part is a reformulation of [4, Th. 5.3.5]. The second part is obtained by contradiction. ∎
@@
We call a random variable ${\bf X}$ in ${\mathbb{R}}^{d}$ isotropic if its distribution is invariant by rotation, i.e., if for any $Q\in O(d)$ and any integrable function $f\colon{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$ we have ${\mathbb{E}}(f({\bf X}))={\mathbb{E}}(f(Q{\bf X}))$. For such a random variable, and all integer $k$, we define the quantities:
$c_{k}({\bf X}):=\frac{2^{k}k!}{(2k)!}\frac{{\mathbb{E}}(X^{2k})}{{\mathbb{E}}(X^{2})^{k}},$
@@
Let ${\bf X}$ be an isotropic random variable in ${\mathbb{R}}^{d}$. Then ${\mathbb{E}}({\bf X}{\bf X}^{T})=\lambda I_{d}$ for some $\lambda\in{\mathbb{R}}$.
@@
The definition of isotropy with $f({\bf x})={\bf x}{\bf x}^{T}$ yields that matrix ${\mathbb{E}}({\bf X}{\bf X}^{T})$ commutes with all orthogonal matrices $Q\in O(d)$, and in particular it commutes with all permutation matrices, a property satisfied only by matrices of the form $\lambda I_{d}$. ∎
@@
${\mathbb{E}}({\bf a}_{1}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})=\begin{cases}c_{k}({\bf X})\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathbb{E}}({\bf a}_{\ell}^{T}{\bf X}\cdot{\bf a}_{r}^{T}{\bf X}),&\text{if $n=2k$,}\\ 0&\text{if $n$ is odd.}\end{cases}$
@@
Let $A=({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n})$ be the matrix containing the vectors ${\bf a}_{i}$ in its columns, and define $Y:=A^{T}{\bf X}$. As $X$ is isotropic, the tensor $T({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n}):={\mathbb{E}}(Y^{1}\cdot Y^{2}\cdots Y^{n})={\mathbb{E}}({\bf a}_{1}^{T}{\bf X}\cdot{\bf a}_{2}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})$ is an isotropic and symmetric tensor, and hence of the form (2) for some real constant $\alpha_{p}=c$. We now compute ${\mathbb{E}}(Y^{\ell}\cdot Y^{r})={\mathbb{E}}({\bf a}_{\ell}^{T}{\bf X}\cdot{\bf a}_{r}^{T}{\bf X})={\mathbb{E}}({\bf a}_{\ell}^{T}{\bf X}{\bf X}^{T}{\bf a}_{r})={\bf a}_{\ell}^{T}{\mathbb{E}}({\bf X}{\bf X}^{T}){\bf a}_{r}=\lambda({\bf a}_{\ell},{\bf a}_{r})$ where we used Lemma 1. Finally, when $n=2k$, choosing ${\bf a}_{j}={\bf a}$, for an arbitrary vector ${\bf a}$ shows that $c=c_{k}({\bf X})$. ∎
Using that the moment generating function of the standard normal distribution is $\mathrm{e}^{1/2t^{2}}$, one can show that ${\mathbb{E}}(X^{2k})=\frac{(2k)!}{2^{k}k!}=\#\operatorname{PP}(2k)$, from which we deduce for the case of multivariate normally distributed vectors with mean zero, $c_{k}({\bf X})=1$. This permits to deduce the proof of Isserlis’ Theorem 1.
@@ 3. Perspectives
(see [10, Theorem 4.1]) Consider a normally distributed vector ${\bf Y}\sim{\mathcal{N}}_{d}({\bf 0},\Sigma)$ in ${\mathbb{R}}^{d}$ with mean zero and covariance matrix $\Sigma=AA^{T}$ with $A=({\bf a}_{1},\ldots,{\bf a}_{d})$. Then, for all tensor $T\colon\otimes^{n}{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, we have
${\mathbb{E}}\big{(}T({\bf Y},{\bf Y},\ldots,{\bf Y})\big{)}=\sum_{p\in\operatorname{PP}(n)}\sum_{(k_{1},k_{2},\ldots,k_{n})\in K_{d}(p)}T({\bf a}_{k_{1}},\ldots,{\bf a}_{k_{n}})$

### v2

@@ 1. Isserlis’s theorem
Let ${\bf Y}=(Y_{1},Y_{2},\ldots,Y_{n})^{T}\sim{\mathcal{N}}_{n}({\bf 0},\Sigma)$ be a multivariate normally distributed vector in ${\mathbb{R}}^{n}$ with mean zero. Then
(1) ${\mathsf{E}}(Y_{1}Y_{2}\cdots Y_{n})=\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathsf{E}}(Y_{\ell}Y_{r}).$
The sum is over all partitions of $n$ into disjoint pairs, denoted $\operatorname{PP}(n)$. For odd $n$, $\operatorname{PP}(n)=\emptyset$ is empty and (1) reduces to $E(Y_{1}Y_{2}\cdots Y_{n})=0$, and we recover that odd moments of normal distributions with mean zero always vanish. For even $n$ with $n=2k$, the number of elements of $\operatorname{PP}(n)$ is exactly #$\#\operatorname{PP}(n)=\frac{(2k)!}{2^{k}k!}.$ For $n=4$, we have #$\#\operatorname{PP}(4)=3$ with
@@
${\mathsf{E}}(Y_{1}Y_{2}Y_{3}Y_{4})={\mathsf{E}}(Y_{1}Y_{2}){\mathsf{E}}(Y_{3}Y_{4})+{\mathsf{E}}(Y_{1}Y_{3}){\mathsf{E}}(Y_{2}Y_{4})+{\mathsf{E}}(Y_{1}Y_{4}){\mathsf{E}}(Y_{2}Y_{3}).$
@@ 2. Isserlis’ Theorem as a corollary of the invariant tensor theorem
We shall derive formula (1) as a consequence of the invariant tensor theorem for isotropic tensors [14, Sec. II.9][1, Appendix I][8, §33.2] stating that the linear space of isotropic tensors is spanned by certain elementary isotropic tensors with two inputs.
@@
The first part is a reformulation of [5, Th. 5.3.5]. The second part is obtained by contradiction. ∎
@@
We call a random variable ${\bf X}$ in ${\mathbb{R}}^{d}$ isotropic if its distribution is invariant by rotation, i.e., if for any $Q\in O(d)$ and any integrable function $f\colon{\mathbb{R}}^{d}\mapsto{\mathbb{R}}$ we have ${\mathsf{E}}(f({\bf X}))={\mathsf{E}}(f(Q{\bf X}))$. For such a random variable, and all integer $k$, we define the quantities:
$c_{k}({\bf X}):=\frac{2^{k}k!}{(2k)!}\frac{{\mathsf{E}}(X^{2k})}{{\mathsf{E}}(X^{2})^{k}},$
@@
Let ${\bf X}$ be an isotropic random variable in ${\mathbb{R}}^{d}$. Then ${\mathsf{E}}({\bf X}{\bf X}^{T})=\lambda I_{d}$ for some $\lambda\in{\mathbb{R}}$.
@@
The definition of isotropy with $f({\bf x})={\bf x}{\bf x}^{T}$ yields that the matrix $\Sigma={\mathsf{E}}({\bf X}{\bf X}^{T})$ commutes with all orthogonal matrices $Q\in O(d)$ a property satisfied only by matrices of the form $\lambda I_{d}$. Indeed, if a matrix $\Sigma$ commutes in particular with all orthogonal symmetries $S$, using that $\Sigma$ leaves invariant the eigenspace $\operatorname{Ker}(S+I_{d})$ of $S$, we obtain $\Sigma{\bf x}\in\operatorname{Vect}({\bf x})$ for all ${\bf x}\in{\mathbb{R}}^{d}$. This later property yields $\Sigma{\bf x}=\lambda{\bf x}$ for all ${\bf x}\in{\mathbb{R}}^{d}$ and for some $\lambda\in{\mathbb{R}}$ which is shown to be independent of ${\bf x}$ by contradiction and this concludes the proof. ∎
@@
${\mathsf{E}}({\bf a}_{1}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})=\begin{cases}c_{k}({\bf X})\sum_{p\in\operatorname{PP}(n)}\prod_{(\ell,r)\in p}{\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}\,{\bf a}_{r}^{T}{\bf X}),&\text{if $n=2k$,}\\ 0&\text{if $n$ is odd.}\end{cases}$
@@
Let $A=({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n})$ be the matrix containing the vectors ${\bf a}_{i}$ in its columns, and define $Y:=A^{T}{\bf X}$. As $X$ is isotropic, the tensor $T({\bf a}_{1},{\bf a}_{2},\ldots,{\bf a}_{n}):={\mathsf{E}}(Y_{1}Y_{2}\cdots Y_{n})={\mathsf{E}}({\bf a}_{1}^{T}{\bf X}\,{\bf a}_{2}^{T}{\bf X}\cdots{\bf a}_{n}^{T}{\bf X})$ is an isotropic and symmetric tensor, and hence of the form (2) for some real constant $\alpha_{p}=c$. We now compute ${\mathsf{E}}(Y_{\ell}Y_{r})={\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}\,{\bf a}_{r}^{T}{\bf X})={\mathsf{E}}({\bf a}_{\ell}^{T}{\bf X}{\bf X}^{T}{\bf a}_{r})={\bf a}_{\ell}^{T}{\mathsf{E}}({\bf X}{\bf X}^{T}){\bf a}_{r}=\lambda({\bf a}_{\ell},{\bf a}_{r})$ where we used Lemma 1. Finally, when $n=2k$, choosing ${\bf a}_{j}={\bf a}$, for an arbitrary vector ${\bf a}$ shows that $c=c_{k}({\bf X})$. ∎
Using that the moment generating function of the standard normal distribution is $\mathrm{e}^{t^{2}/2}$, one can show that ${\mathsf{E}}(X^{2k})=\frac{(2k)!}{2^{k}k!}=\#\operatorname{PP}(2k)$, from which we deduce for the case of multivariate normally distributed vectors with mean zero, $c_{k}({\bf X})=1$. This permits to deduce the proof of Isserlis’ Theorem 1.
@@ 3. Perspectives
(see [11, Theorem 4.1]) Consider a normally distributed vector ${\bf Y}\sim{\mathcal{N}}_{d}({\bf 0},\Sigma)$ in ${\mathbb{R}}^{d}$ with mean zero and covariance matrix $\Sigma=AA^{T}$ with $A=({\bf a}_{1},\ldots,{\bf a}_{d})$. Then, for all tensor $T\colon\otimes^{n}{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}$, we have
${\mathsf{E}}\big{(}T({\bf Y},{\bf Y},\ldots,{\bf Y})\big{)}=\sum_{p\in\operatorname{PP}(n)}\sum_{(k_{1},k_{2},\ldots,k_{n})\in K_{d}(p)}T({\bf a}_{k_{1}},\ldots,{\bf a}_{k_{n}})$